At the dog park their are several dogs with their owners. Counting heads, there are 12, counting legs, there are 38. How many dogs and owners are there?
Answer by stanbon(75887)

At the dog park their are several dogs with their owners. Counting heads, there are 12, counting legs, there are 38. How many dogs and owners are there?
----------------
Equations:
d + w = 12
4d+2w = 38
----
Multiply thru 1st equation by 2:
2d + 2w = 24
4d + 2w = 38
---------------
Subtract and solve for "d":
2d = 14
d = 7 (# of dogs)
----
Solve for "w":
d + w = 12
7 + w = 12
w = 5 (# of owners)
=========================
Cheers,
Stan H.
=================